Calculate the mole fraction of each component of a solution containing 65g of ethanol (C2H6O) in 350g water.


1
Expert's answer
2021-02-24T03:37:50-0500

Moler mass of Ethanol ("C_2H_6O" )=46 g/mol

so moles of ethanol="\dfrac{}{}" "\dfrac{65}{46}" =1.41 moles


Now moles of water = "\dfrac{350}{18}" =19.45 moles


Total moles = 1.41+19.45=20.86 moles


Mole fraction of ethanol= moles of ethnol/total moles

"\Rightarrow" "\dfrac{1.41}{20.86}" =0.067


Mole fraction of water=moles of water/total moles

"\Rightarrow" "\dfrac{19.45}{20.86}" =0.932